OpenStudy (helpplease456711):

HELP WILL MEDAL AND FAN There is a traiffic light at the intersection of Walnut and Lawrence avenue. The traffic light on Walnut street follows a cycle. It is green for 50 seconds, yellow for 10 seconds, and red for 30 seconds. As you travel along Walnut and approach the intersection, what is the probability that the first color you see is red

OpenStudy (helpplease456711):

@TheSmartOne

OpenStudy (helpplease456711):

I think it's 1/3 I'm not sure

jimthompson5910 (jim_thompson5910):

Yes you are correct. 30 of the 90 seconds (50+10+30=90) is when the light is red 30/90 = 1/3

OpenStudy (helpplease456711):

Thank you

jimthompson5910 (jim_thompson5910):

no problem